If 3x - 4y = 5, then what is the value of y/3 - x/4?
Answer Choices:
a) 5
b) -5/12
c) 1/5
d) -1/5

Answer by Boreal(15235)

4y=3x-5
y=(3/4)x-(5/4)
y/3=(1/4)x-(5/12)
y/3-(x/4)=-5/12
B